OpenStudy (anonymous):

Hi guys. If been find it hard to solve the following equation: 16 + 8 Times Square root of 5 all of the above divided by 27 and then all of the equation square rooted by 3). Could you guys please show me the method. Thx

OpenStudy (anonymous):

Yes all of the equation square rooted by 3

OpenStudy (anonymous):

So you want \[ \Large \sqrt[3]{\frac{16+8\sqrt{5}}{2}} \]

OpenStudy (anonymous):

\[3\sqrt{\frac{ 16+8\sqrt{5} }{ 27 }}\]

OpenStudy (anonymous):

Yes that's the one